OpenStudy (anonymous):

A cone has a base radius of 2 millimeters and a height of 24 millimeters. What is the volume of the cone? A. 384π mm3 B. 96π mm3 C. 16π mm3 D. 32π mm3

OpenStudy (anonymous):

i got 100.4 but that is not one of the answer

OpenStudy (anonymous):

i dont know what I am getting wrong

OpenStudy (anonymous):

@isaidavila

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

Volume of a cone: \[V=\pi*r^2*\frac{h}{3}\]

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

i am still getting 100.48

OpenStudy (anonymous):

type find the volume of a cone in google and theres a thing where you type in the info and gives the answer lol

OpenStudy (anonymous):

2^2 * (24/3) = 4*8 = 32pi

OpenStudy (kropot72):

But you need to include the symbol for pi in your answer.

OpenStudy (anonymous):

i did and pi equals to 3.14

OpenStudy (anonymous):

so the answer is 32

OpenStudy (anonymous):

THANK YOU WHO EVER HELPED ME